### v1

# Microstructural characterization to reveal evidence of shock deformation in a Campo del Cielo meteorite fragment

Abstract
The study of meteorites and their microstructures is a topic which spans multiple fields of research, such as earth & planetary sciences, meteoritics, geosciences, and materials science. For materials scientists and engineers, the extreme and unusual conditions which these microstructures form allow for insight into materials which would exist at the edge of our thermomechanical processing abilities. One such microstructure found in low-shock event iron meteorites is Neumann bands. These bands are an array of lenticular deformation twins that form throughout the Fe-Ni matrix with numerous intersections, resulting in many high stress and strain regions within the material’s surface. The existence of these regions and the shocks that formed them encourage atypical strain accommodating mechanisms and structural changes of the material. However, direct investigation of the deformation twin intersections and the microstructural behaviour in and around these regions has been limited. In this work, investigation of these regions in a Campo del Cielo meteorite fragment, with electron backscatter diffraction (EBSD) and forescatter electron (FSE) imaging, revealed two primary findings: high-intensity pattern doubling mirrored across the {110} band at twin-twin intersection and microband formation across the sample surface, which suggest multilayer twinning and constraint of the crystal structure at points of twin-twin intersection. Microbands were found to form along the {110} plane and in regions near Neumann bands. The simultaneous existence of Neumann bands (microtwins) and microbands is presented here for a BCC material, and it is believed the Neumann band and microbands formed during different types and/or shock events from one another. The presence of both Neumann bands and microbands within a BCC iron meteorite is previously unreported and may be valuable in furthering our understanding of shock deformation within iron-based materials.
1 Introduction
Meteorites and their fragments provide insight into the formation of materials on earth and in space. In brief, a meteorite is a meteor that survives its passage through the atmosphere of the earth and strikes the ground. These meteorites often contain very interesting microstructures that are uncommonly found on earth, as they have unique chemistry and thermomechanical processing histories (e.g. very long cooling periods) that depend on the meteor’s formation, movement through space, and location of impact. When the meteor becomes a meteorite, it may experience short-lived rapid shock which can result in some unusual characteristic microstructural features as deformation occurs at very high strain rates (estimated between and ) [1], often with internal aftershocks. Therefore, the ultimate microstructure of a meteorite fragment is dependent on the impact and thermal histories of the host meteor and the resulting meteorite [2, 3], in a similar manner to how the microstructure of an engineering component is influenced by the manufacturing process. These high strain rate microstructures, resulting from impact shocks, can be characterized to understand these otherworldly materials, and provide insight into the extremes of our accessible thermomechanical processing envelope.
The present manuscript focuses on the Campo del Cielo meteorite, which is an iron-rich Fe-Ni metallic meteorite. Major phases in the Fe-Ni system include kamacite () and taenite (). The kamacite phase is closely related to the ferrite phase in steel, with a body centred cubic (BCC) crystal structure and a chemistry that is Fe-rich with typically 3.3-5.8 wt% Ni [4, 5]. The taenite phase is closely related to the austenite phase in steel, with a face centered cubic (FCC) crystal structure and a chemistry that is Fe-rich with typically 31 to 54 wt% Ni [4, 5]. Additionally, the presence of other elements, such as P, S, and Co, results in the precipitation of different minerals within the iron matrix. A common mineral precipitate that is found is schreibersite (also known as rhabdite), an iron-nickel phosphide mineral with a tetragonal crystal structure and appears as a blocky precipitate. This precipitate is found to form in preferential orientations, with the orientation relationship given in Eqn. 1 and 2 [6]. Iron meteorites are also known to have other microstructural features which form in preferential orientations, such as Neumann bands.
Neumann bands are an observed form of deformation twinning, where the twins are found within the kamacite phase along the {112} planes. The formation of these bands is a direct result of the shock incurred from the meteorite’s impact [3], with each twin providing a high rate strain accommodation mechanism that will be formed as one of twelve possible twin variants with respect to the parent BCC-grain [7, 8]. For these high rate deformation twins in BCC-kamacite, it is thought that the nuclei of these twins first form due to an accumulation of stacking faults along {112} planes [9], with twin propagation initiating once a critical shear stress value is reached [10]. The twins are thought to form as an oblate spheroid, which present within a 2D planar section as elongated ovoids.
Neumann bands tend to extend (at least) across the surface of iron meteorites and exist in various thicknesses, lengths and crystal orientations. The twins tend to form in groups of similar direction and orientation, but they do not all form simultaneously. Rather, families of Neumann bands form as a result of different shock waves that move through the meteorite and it is likely that each band that extends in the same direction within a large single crystal matrix is related to the individual variant of the deformation twin. When multiple families of bands form, this results in twin growth and intersections through one another.
The growth and intersection of these bands in kamacite, and other BCC metals, are dependent on the twinning, slip and shear planes within the metal-matrix grain. After the initial formation of a band, it will undergo forward growth along the twin shear direction, {112}, through deformation twinning. As the twin grows it generates highly strained regions within the matrix, thereby raising the driving stress needed for further propagation [11, 12, 13]. This increase in required driving stress leads to dislocation mobility and slip being a more energetically favourable deformation mechanism. Resulting in outward growth and thickening of the Neumann bands until forward growth becomes energetically favourable again.
Single slip deformation in BCC metals is also known to result in the formation of another visually unique microstructural feature called microbands [14, 15]. Microbands appear as linear bands, coincident with the {110} planes in BCC metals [16], and form initially from strain-induced primary slip planes with independent dislocations. Shear stress acting on these slip planes then results in a polarized arrangement of the dislocations and the formation of double dislocation walls. A more in-depth explanation of the microband formation mechanism is given by Huang and Gray [15]. While these bands are mechanistically independent from Neumann bands, they may share traces of {110} with Neumann band traces of {112} along the [110] orientation [16]. With that said, microbands and Neumann bands have been previously observed within iron and steel materials separately [17, 18] and in lab-based impact tests in stainless steel "crater-related microstructures are dominated by deformation microtwins with a few intermixed microbands" [19]. However, we are not aware of previously documented simultaneous existence of these microstructural features for iron meteorites.
The Neumann bands, their twelve possible formation directions and their frequent intergrowth result in a complex microstructure which contains twin intersections with high strain gradients. These intersections and the regions surrounding them are of high interest for investigation using Electron Backscatter Diffraction (EBSD). Previous works have used this technique to investigate iron meteorites [3, 20, 21]. However, there has been limited study of the intersections of the deformation twins and the microstructural behaviour in these regions.
The present work aims to remedy this by further investigating the high stress and strain areas created by the Neumann bands and their intersections present within the microstructure of iron-based meteorites. In particular, abnormal material behaviour and phenomena are to be investigated using EBSD. Similar works have been performed previously for both titanium deformation twins [22] and for other iron-based materials [10]. However, there is a lack of investigation into these areas for the Campo del Cielo meteorite group.
2 Materials and Methods
2.1 Material Preparation
A meteorite fragment from the Campo del Cielo meteorite group (purchased from Rocks Gems Canada, Whistler, BC) was metallographically polished using 400, 600, 800, and 1200 grit silicon carbide for 10 minutes each, with the sample rotated after 5 minutes per grit. This was followed by polishing with a and diamond suspension, and a colloidal silica suspension for 6 minutes each (again rotating by halfway through). Lastly, broad ion beam (Ar-ions) polishing was performed using a Gatan PECS II system for 1 hour at tilt, , 1 rpm, and no modulation.
2.2 Material Analysis
Electron microscopy was performed on a Tescan AMBER-X. Secondary Electron (SE) images were obtained using a beam current, accelerating voltage, and a working distance of 13-14 . EBSD was performed with an Oxford Instruments Symmetry S2 detector using beam current and accelerating voltage, with the sample loaded into a pre-tilted holder. EBSD acquisition was performed three times with increased magnification on regions of interest (ROI). The first ROI was imaged with an acquisition speed of , a resolution of 156x, and an exposure time of , the second ROI at , 1244x, and and the third ROI at , 622x, and . Further Forescatter electron (FSE) imaging was performed with an increased dwell time of at a working distance of , , and .
Post-acquisition analysis of the EBSD data was performed using Aztec Crystal, using MapSweeper which performs a ’template matching’ approach of comparing each captured diffraction pattern with a dynamical-diffraction based simulation and this improves the indexing precision [23], and subsequent analysis within MTEX 5.10.2 [24] in MATLAB. For one map (shown in Figure 3), high (angular) resolution EBSD (HR-EBSD) analysis was performed using CrossCourt Rapide v4.6 to perform total GND analysis in the BCC-phase assuming <110> dislocation types, using remapping based cross correlation with 40 ROIs and [25].
The sample was mounted flat for energy dispersive X-ray spectroscopy (EDS/EDX), and area, point and mapping was performed using an Oxford Instruments EDS analysis using an Ultim Max 170 detector. Analysis was performed at , , and a working distance of , using Oxford Instruments Aztec with the Pulse Pile up correction.
3 Results
Fig. 1 shows an overview of the area studied, and reveals that the meteorite fragment was found to be composed of kamacite with a large number of cuboid schreibersite precipitates, as confirmed with EBSD (crystallography) and EDS (chemistry). EDS results were as expected for kamacite and schreibersite. The kamacite matrix was composed of 88.8-94.0 wt% Fe and 5.8-6.0 wt% Ni, any remainder was attributed to be carbon. The schreibersite precipitates had elemental composition of 42.6-45.2 wt% Fe, 35.0-36.5 wt% Ni, 13.6-12.3 wt% P, and 8.8-5.9 wt% C. The map sum elemental composition of the meteorite fragment was 88.2 wt% Fe, 5.9 wt% Ni, 5.8 wt% C, 0.1 wt% P. The carbon map is not provided in Fig. 1 as it was found dispersed across the sample and may be from additional surface contamination.
Neumann bands were seen to be interlaced over the entire surface with crystal orientations being consistent amongst parallel bands. The boundaries of these bands were found to be <111> twins with a misorientation. Areas of elevated lattice rotation gradient (i.e. strain gradient) can be seen around the intersection point of the Neumann bands, shown by the slight changes of orientation in the IPF plots. These regions are most visible in the IPF-X of Fig. 1 and further investigation was performed at the isolated twin tip region in Fig. 2.
A greater magnitude of misorientation was seen in the lower twin present in Fig. 2. The IPF maps showed minimal orientation changes in the kamacite surrounding the isolated twin tip, however an increase in misorientation is seen extending outward from the twin tip. The kernel average misorientation (KAM) was seen to be largest in this region of the tip, though the magnitude was low (approx. ).
The isolated twin was found to have increased misorientation at a second location near its middle. The source of misorientation near the middle appears to be the initiation of a new twin which had not fully formed into a Neumann band. However, the embryo of a Neumann band can be seen through the KAM in Fig. 2. The presence of these embryos is common on the sample surface and are likely indicative of the direction which a new Neumann band would form with additional shock.
Further analysis was performed using Weighted Burgers Vector (WBV) [26] to analyze the dislocations present within the isolated tip ROI (Fig. 2). The WBV magnitude reveals a greater number of dislocations within the tip front region that fan out and upward toward the twin band present in the upper right. A large WBV magnitude is observed found near the middle of the twin, of course this is also revealed as regions of measured KAM (which is related to the GND density and the WBV analysis). The presence of the potential Neumann band formation embryo was similarly visible using WBV magnitude compared to the KAM. These dislocation structures are most evident within the WBV-Z maps.
A further region of interest was identified near the isolated twin tip, shown in Fig. 3. The region in question was identified to have a high KAM value and a large number of dislocations within a small twin intersection. Initial indexing of this location failed and direct investigation resulted in the discovery of the Electron Backscatter Patterns (EBSPs) seen in Fig. 3.
SE imaging also lead to the discovery of a seemingly ‘rippled’ surface as seen in Fig. 4. These ‘ripples’ were seen to run parallel to one another, with a consistent spacing of approximately . The ripples were found to be present over large distances and were shifted when near Neumann bands.
A high magnification analysis of the ‘ripples’ was performed and is provided in Fig. 5. The ‘ripples’ were seen to extend parallel to Neumann bands present in the material surface but did not appear to change direction when crossing large Neumann bands. Pole figures were created for the {110}, and {112} planes and [111] direction. Surface traces indicated that the ‘wavefront’ (red-blue and blue-blue traces, Fig. 5) of these ripples aligned with {110}. The ripple-ripple interface was found to align with {112} (green-blue trace, Fig. 5).
4 Discussion
This Campo del Cielo fragment was found to be covered in Neumann bands, with numerous intersections points between these bands. Closer investigation of isolated Neumann bands/twins revealed a network of dislocations (Fig. 2) from an isolated twin tip to the Neumann band twin present above it and a higher WBV magnitude, and therefore dislocation density, along the boundary of the lower Neumann band.
The dislocation density on the lower Neumann band varied along the boundary, with areas of high density dotted along the upper side of the twin grain boundary. It is noted that the increased density is not seen in the IPF-Y and Z (Fig. 2) on the lower side of the grain boundary. This may be a result of the grain growth mechanism for BCC deformation twins. As found by Zhong et al. for BCC tantalum, the twin growth in the BCC metals may occur through a nucleation-controlled mechanism or a growth-controlled mechanism [11]. The later mechanism occurs with the twin grain having a coherent twin boundary (CTB) on one side and an incoherent twin boundary on the other. The widening of the twin occurs on the CTB side and results in a gradual build up of lattice strains and strain energies on the incoherent side until the energy required for nucleation of another CTB to occur is overcome.
This growth-controlled mechanism is believed to be present in the lower Neumann band, with the CTB being the lower side of the band and the incoherent boundary being the upper side. High strain energies and raised dislocation density are seen on the upper side and indicates that the twin grain has not reached the critical strain energy required to initiate further growth. From Fig. 2 it can also be seen that these high strain regions are present at locations when the band is thinner, but adjacent to thicker regions. While the growth occurs rapidly once the energy barrier is overcome [11], it may be that when this barrier is reached that it is not done uniformly across the twin grain boundary.
The isolated twin tip with the network of dislocations does not exhibit this same behaviour. Rather, the twin tip does not have a CTB and incoherent boundary but instead exhibited a lower KAM and WBV magnitude along its body, with increased magnitudes at and near the tip. It is possible that the increased dislocation density near the tip is a precursor to new grain boundary formation through a nucleation-controlled mechanism. Similar behaviour has been simulated previously by Sainath and Nagesha for the nucleation and initiation of a BCC iron twin [27].
However, if this twin tip were to further propagate it would be outward from the tip point, rather than upward toward the Neumann band in the upper right of the WBV magnitude in Fig. 2.
Further analysis of the smallest region of interest lead to the discovery of the unusual EBSP shown in Fig. 3 in this region of very large deformation near the edge of the twin-twin interface. The twin and matrix EBSPs show that there is a common {110} band within these diffraction patterns, and the unindexed region (in region 3 of the map) has a pattern which contains a single diffraction band superimposed upon a noisy ’background’ (which has some evidence of other diffracting features). In general, in the presence of a strain gradient within the interaction volume, it is well known and well understood that the EBSD-pattern can blur and become harder to index. In this unusual example, all but one band has blurred, and this indicates that the strain gradient for this volume is highly restricted and there is crystalline material with a (relatively) undisturbed {110} plane and one could imagine that this is the extreme case of ’pattern doubling’ that could be related to either a very high strain gradient or multilayer twinning [28] in this localized region. This motivated the HR-EBSD based GND analysis, which supports the idea that there is a localized strain gradient in this location due to the presence of a very high GND density near this region.
A multilayer twin is a section of the material where multiple twin deformations are stacked atop one another. These multilayer twins are denoted by the number of packing layers between the twins, with density of twin increasing as the number is lower [28]. For example, a 9-layer twin has 9 layers (ABCABCABC) between it and the twin deformation plane (illustration provided in [28]). When the electron beam direction is parallel to the twin plane of this stack it results in overlap of EBSP of each twin, causing constructive interference near the twin plane band and increasingly destructive interference further away [28]. This behaviour has been simulated previously by Li et al. for FCC iron down to 3-layer twins. The simulated EBSP presented by Li et al. appears similar in nature to the unusual EBSP seen in this work, however the EBSP seen here experiences greater amounts of interference (constructive and destructive). Therefore it is hypothesized that if multilayer twinning is present in this EBSP, then it is at a greater degree than 3-layer.
Constraint of the crystal structure may also play a part in the high {110} presence seen in the EBSP. Intersection of the two twin bands results in the generation of shear stress and movement through specific slip systems. Those being {110}, {112}, and {123} for BCC metals, in order of increasing rarity [29]. The most common slip plane, {110}, is known to be related to the formation of Neumann bands, with {110} acting as a twin plane for the formation of Neumann bands [7]. It is suggested that the twin-twin intersection and the shear stress involved results in slip along the {110} plane, constraining the microstructure and its dislocations. Thereby, causing a high intensity {110} band to be seen at the site of exit for the twin-twin intersection.
SE imaging of this region also showed an interesting surface phenomena. A ripple-like appearance was seen (Fig. 5) that seemed to travel along planes similar to the large Neumann bands. Surface traces on the {110} and {112} pole figures were performed, given the known relation of these planes to Neumann bands [8, 7]. This revealed an overlap with the {110} plane and ‘wavefront’ of the ripple, and for the {112} plane with the ripple-ripple interface. Additional ripples were seen over large regions of the sample, shown in Fig. 4, that have the same ‘wavefront’, but are not seen to have the same ripple-ripple interfaces as in Fig. 5.
The relation seen for the ‘wavefront’, or walls, of these ‘ripples’ with {110} and their presence in and around high strain regions suggests that these may be microbands. The development of microbands in BCC-iron and other BCC materials is a known phenomena induced through rapid shock [30, 14], which forms along BCC slip systems ({110}, {112}) [30, 31]. Similar behaviour is found here with the ‘wavefront’ coinciding with the preferential slip plane of microband formation. The ripple-ripple interface seen along {112} is believed to be a result of deformation twin growth, resulting in shear deformation of the surrounding matrix. An illustration is provided in Fig. 5 e) to show the relationship of the ripples to the slip direction, shear direction, and the microbands.
This would indicate that the ripples and Neumann bands did not form simultaneously. The overlapping of these ripples with the Neumann bands in Fig. 4 and lack thereof in Fig. 5 a) further reinforces this. The lack of overlap seen in the latter is indicative that the deformation twin formed after the ripples, likely due to a different shock event. Different types of shock has been found in the past to affect what microstructural features will develop in FCC materials, with plane shock and spherical shock inducing deformation twinning and microband formation, respectively [32]. While this relationship was developed using FCC materials, it has been proposed by Huang and Grey that the formation mechanism of microbands may be similar in FCC and BCC materials [15]. Rather, the specific material dictates the propensity for strain accommodation to occur through microband formation or deformation twinning.
5 Conclusion
The Campo del Cielo meteorite fragment was found to be composed of a kamacite matrix with blocky schreibersite precipitates present throughout. Deformation twin grains, called Neumann bands, were found to cover the surface of the sample with a high amount of intersection. These Neumann bands and their intersections resulted in high stress and strain regions within the sample surface that were detectable through the use of EBSD-based analysis.
An investigation of isolated Neumann bands using WBV analysis revealed a network of dislocations extending outward from an isolated twin tip and spots of high dislocation density along one side of a Neumann band. The high dislocation density on a single twin boundary was believed to be due to a growth-controlled propagation mechanism dominating for this grain. The high dislocation density and elevated strain energies were found to correspond to incoherent twin boundary side of the Neumann band and may be a result of non-uniform growth on the incoherent twin boundary. Additionally, networks of dislocations were seen to extend outward from the isolated twin tip and are believed to be indicative of further twin propagation through a nucleation-controlled growth mechanism.
Closer investigation was performed into a high strain region of the isolated Neumann bands where a twin-twin intersection was discovered to be the source. Unusual EBSPs were found at the exit point of this intersection with a high intensity {110} band with blurry and lower intensity bands being mirrored across the {110} band. It is suggested this phenomenon was a result of multilayer twinning at the exit point leading to constructive interference for the {110} band and constraint of the crystal structure to the {110} plane due to dislocation slip through the twin plane of the larger Neumann band.
Additionally, a unique surface phenomenon was seen around the Neumann bands upon high magnification imaging at multiple locations of the sample. A rippled microstructural feature was found surrounding the Neumann bands, or overlapping with them in other instances. These ripples were all seen to have a ‘wavefront’ which coincided with the {110} plane of kamacite, and some with a ripple-ripple interface coinciding with the {112} plane of kamacite. These microstructural features are suggested to be a microband strain accommodating feature within the iron meteorite which are formed through slip deformation. Furthermore, it is suggested that the formation of the Neumann bands and the ripples occurred due to separate shock events, planar and spherical shock respectively.
To summarize, EBSD-based investigation of high strain Neumann band regions in the Campo del Cielo meteorite fragment revealed two highly unusual phenomena, the ‘ripple’ microbands across the surface and the mirrored high intensity {110} band EBSP present at a Neumann band intersection. The presence of such phenomena may be valuable in furthering our understanding of high strain rate microstructural changes within both terrestrial and interstellar iron-based materials.
Acknowledgements
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) [Discovery grant: RGPIN-2022–04762, ‘Advances in Data Driven Quantitative Materials Characterisation’]. The AMBER-X pFIB-SEM was funded by the BCKDF and CFI-IF (#39798: AM+).